# Update-Aware Robust Optimal Model Predictive Control for Nonlinear Systems

### v1

@@ Abstract
Index Terms:
Predictive control for nonlinear systems, Robust control, Optimal control.
@@ II Problem Description
Let $f_{k}(\cdot,\cdot,\cdot,\cdot):$ $\mathcal{X}\times\mathcal{X}\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}^{n_{x}}$ be the function describing the discrete-time dynamics of the system with state $x_{k}$in $\mathcal{X}$subseteq $\mathbb{R}^{n_{x}}$ at time-step $k$, associated control inputs $u_{k}$in $\mathcal{U}$subseteq $\mathbb{R}^{n_{u}}$, disturbances $w_{k}$in $\mathcal{W}$subseteq $\mathbb{R}^{n_{w}}$, such that
@@
We consider the problem of controlling this system between time step 00 and a known final time step $N$in $\mathbb{N}$. We denote by $x\in\bar{\mathcal{X}}\coloneq\mathcal{X}^{N+1}$ the collection of states from $k=0$ to $k=N$, by $u\in\bar{\mathcal{U}}\coloneq\mathcal{U}^{N}$ the control inputs between 00 and $N-1$, and by $w\in\bar{\mathcal{W}}\coloneq\mathcal{W}^{N}$ the uncertainty sequence between 00 and $N-1$. The dynamics of (1) between 00 and $N$ can then be combined with an initial condition constraint on $x_{0}$ into a single function describing the evolution of the state,
@@
$\mathcal{U}^{*}\coloneqq\left\{u\left|\>\forall{w}\in\bar{\mathcal{W}},\,\forall{x}\in X(u,w):\>\>g(x,u,w)\leq 0\right.\right\}.$ (6)
@@
In [16], we showed how the problem of (7) could be cast as a SIP and an open-loop solution could be obtained through the application of a local reduction method, as described in Section IV. Here, we examine the case where $u$ need not be computed in its entirety at time step 0, and show how updating $u$ within a decreasing horizon context can lead to performance improvements. Through the rest of this paper, we assume that the control inputs can be re-computed at every time step between 00 and $N$.
In order to describe our approach, we now introduce $x_{k}^{+}\in\mathcal{X}_{k}^{+}\coloneqq\mathcal{X}^{N-k}$ to refer to the components of $x$ corresponding to the time steps between $k+1$ and $N$, $u_{k}^{+}\in\mathcal{U}_{k}^{+}\coloneq\mathcal{U}^{N-k}$ to refer to the decisions of $u$ that are to be made between $k$ and $N-1$. Similarly, we define $w_{k}^{+}\in\mathcal{W}_{k}^{+}\coloneq\mathcal{W}^{N-k}$ to refer to the time-varying disturbances of $w$ that are realized between $k$ and $N$. We also use the notation $(\cdot)_{k}^{-}$ to refer to any quantity between 00 and $k-1$, with the exception of $x_{k}^{-}$, which includes $x_{0}$ through $x_{k}$.
@@
$J_{k}^{*}(z_{k},u_{k}^{+})\coloneqq\max_{\begin{subarray}{c}\vphantom{\sum}w_{k}^{+}\in\,\mathcal{W}_{k}^{+},\\[1.0pt] x_{k}^{+}\in\,\mathcal{X}_{k}^{+}(z_{k},u_{k}^{+},w_{k}^{+})\end{subarray}}J_{k}(z_{k},x_{k}^{+},u_{k}^{+},w_{k}^{+}).$ (11)
@@
Meanwhile, if $k<N-1$ and we know that the controller will be updated in the future, we only need to compute the next segment in the controller trajectory $u^{*}_{k}(z_{k})$, the control input that, in the worst-case realization of $w_{k}$, leaves the system in the best possible position to minimize the total cost by the time the control input can next be updated at $k+1$. $u^{*}_{k}(z_{k})$ also needs to ensure that future constraints remain satisfiable for all possible uncertainties. The rest of the control trajectory, meanwhile, cannot be computed at time $k$ since it relies on information that will only be available in the future. In Section III, we show how to overcome this limitation to propose a novel update-aware robust optimal MPC algorithm that accounts for the acquisition of future information in its decision making process.
@@ III-A Algorithm Description
$\begin{split}&\mathcal{U}^{*}_{k}(z_{k})\coloneqq\\[3.0pt] &\left\{v_{k}\in\mathcal{U}\left|\begin{array}[]{l}\forall{\omega}_{k}\in\mathcal{W},\\[1.0pt] \forall{s}_{k}\in X_{k}(z_{k},v_{k},\omega_{k})\end{array}:\begin{array}[]{l}g_{k}(z_{k},s_{k},v_{k},\omega_{k})\leq 0\\[1.0pt] \zeta_{k+1}=(z_{k},s_{k},v_{k},\omega_{k})\\[1.0pt] \exists\,v_{k+1}\in\mathcal{U}^{*}_{k+1}(\zeta_{k+1})\end{array}\right.\right\}\end{split}$ (15)
@@ III-B Theoretical Guarantees
Therefore, $\mathcal{U}_{k}^{*}(z_{k})$ and $\bar{\mathcal{U}}_{k}^{*}(z_{k})$ only differ by the ordering of the $\forall$ and $\exists$ operators. Then by universal instantiation, a member of $\bar{\mathcal{U}}_{k}^{*}(z_{k})$ must also be a member of ${\mathcal{U}}_{k}^{*}(z_{k})$, since $\exists A_{1}\forall A_{2}\implies\forall A_{2}\exists A_{1}$ given any $A_{1}$ and $A_{2}$. Members of ${\mathcal{U}}_{k}^{*}(z_{k})$, meanwhile, do not necessarily belong to $\bar{\mathcal{U}}_{k}^{*}(z_{k})$.
@@
$\begin{split}&\bar{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})\coloneqq\left\{v_{k}\in\mathcal{U}\left|\,\exists v_{k+1}\in\mathcal{U}\left|\,\ldots\left|\,\exists v_{N}\in\mathcal{U}\left|\vphantom{\begin{array}[]{c}\\[1.0pt] \end{array}}\right.\right.\right.\right.\right.\\[1.0pt] &\left.\begin{array}[]{l}\forall{\omega}_{k}^{+}\in\mathcal{W}_{k}^{+},\\[1.0pt] \forall{s}_{k}^{+}\in X_{k}^{+}(z_{k},v_{k}^{+},\omega_{k}^{+})\end{array}:\right.\left.\>J_{k}(z_{k},s_{k}^{+},v_{k}^{+},\omega_{k}^{+})\leq\gamma_{k}\vphantom{\begin{array}[]{c}\\[1.0pt] \end{array}}\;\right\}.\end{split}$ (20)
Using the same logic as in Theorem 1, we know that $\bar{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})\subseteq{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})$ $\forall z_{k}\in\mathcal{Z}_{k},\gamma_{k}\in\mathbb{R}$. It then follows that $\left\{\bar{\mathcal{U}}_{k}^{*}(z_{k})\,\cap\,\bar{\mathcal{J}}^{*}(z_{k},\delta_{k})\right\}\subseteq\left\{{\mathcal{U}}_{k}^{*}(z_{k})\,\cap\,{\mathcal{J}}^{*}(z_{k},\delta_{k})\right\}$. Therefore, the feasible space for the problem of (19) is contained in that of (17b), and given that the two objectives are the same, we have that $\gamma_{k}(z_{1})\leq\bar{\gamma_{k}}(z_{k})$ $\forall z_{k}\in\mathcal{Z}_{k}$.
@@
Proof 3
Recursive feasibility follows trivially from (15) since $u^{*}_{k}(z_{k})\in\mathcal{U}_{k}^{*}(z_{k})\implies u_{k+1}^{*}(z_{k+1})\in\mathcal{U}_{k+1}^{*}(z_{k+1})$ $\forall z_{k}\in\mathcal{Z}_{k}$, $w_{k}\in\mathcal{W}$, where $u_{k}=u_{k}^{*}(z_{k})$, $x_{k}=X_{k}(z_{k},u_{k},w_{k})$, and $z_{k+1}=(z_{k},x_{k},u_{k},w_{k})$.
@@
Proof 4
This follows directly from the knowledge that $u_{k}\in\mathcal{U}^{*}(z_{k})\,\cap\,\mathcal{J}^{*}(z_{k},\gamma^{*}_{k})$ under (17a). After applying $u_{k}^{*}(z_{k})$, $\mathcal{J}_{k+1}^{*}(z_{k+1},\gamma_{k}^{*})$ must be non-empty from (16). Similarly, $\mathcal{U}^{*}_{k+1}(z_{k+1})$ is non-empty from Theorem 3. Therefore, $\gamma_{k}^{*}(z_{k})$ is guaranteed to be in the solution space of (17b) for $z_{k+1}$ and the optimal solution can only improve on the previous time-step so that $\gamma^{*}_{k+1}(z_{k+1})\leq\gamma^{*}_{k}(z_{k})$.
@@ IV Solution by Local Reduction
While the problem of (17a) gives us the optimal continuation to our control input trajectory, it remains challenging to solve numerically. We address this by noticing that (17a) is a nested sequence of $N-k+1$ existence-constrained robust optimal control problems, which, under Assumptions 14 can be rewritten as a nested sequence of existence-constrained SIPs using the approach of [16]. Then, from [16], (17) can be reformulated into the sequence of problems
@@
where $\tilde{P}_{k}(z_{k})\coloneqq P_{k}(z_{k})$ if $P_{k}(z_{k})$ is feasible and $\infty$- ∞ if not, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are members of a simplex $\Lambda$, and $\epsilon$ is the smallest positive constant. Substituting (27) for (24b) in (24) then finalizes the conversion into a regular SIP. Notice that solving the constraint generation problem for (27) requires solving $P_{k}^{k+1}(\cdot)$, preserving the recursive structure.
@@ V Results
In order to test the performance of our proposed update-aware robust optimal MPC algorithm, we compare it to a more standard robust optimal control scheme that applies the control of (12a) instead of (17a) on a simulation of a planar quadrotor. The model has continuous-time states $[r,\dot{r},s,\dot{s},\psi,\dot{\psi}]$, where $r$ is the quadrotor’s horizontal position, $s$ is the quadrotor’s height, and $\psi$ is the quadrotor’s tilt angle, as illustrated in Figure 1.
@@
We then obtain the discrete-time dynamics at time step $k$ through the application of a midpoint rule with sampling time $T_{s}=0.5$, and where $x^{1}_{k}$ through $x^{6}_{k}$ are the discretized states at time $k$ corresponding to the continuous-time values $[r,\dot{r},s,\dot{s},\psi,\dot{\psi}]$ and the discretization is such that $k=0$ maps to $t=0$. We examine the behavior of the simulation over 6 time steps, with $\mathcal{K}=\{0,1,2,3,4\}$.
The system is initialized from the origin with the state $x_{0}=[0,0,0,0,0,0]$. We also require that the system’s $x^{1}$ coordinate remain between $-c$ and $c$ at all times and that its $x^{3}$ coordinate remain positive by satisfying the inequalities
@@
$\mathcal{U}=\left\{u\left|\begin{array}[]{r}-2\leq u^{1}_{k}\leq 2\\ -2\leq u^{2}_{k}\leq 2\end{array}\;\forall k\in\{0,\ldots,4\}\right.\right\}$ (34)
@@
$\mathcal{W}=\left\{w\left|-w_{\text{max}}\leq w_{k}\leq w_{\text{max}}\;\forall k\in\{0,\ldots,4\}\right.\right\}.$ (35)

### v2

@@ II Problem Description
Let $f_{k}(\cdot,\cdot,\cdot,\cdot):$ $\mathcal{X}\times\mathcal{X}\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}^{n_{x}}$ be the function describing the discrete-time dynamics of the system with state $x_{k}$ $\in$ $\mathcal{X}$ $\subseteq$ $\mathbb{R}^{n_{x}}$ at time-step $k$, associated control inputs $u_{k}$ $\in$ $\mathcal{U}$ $\subseteq$ $\mathbb{R}^{n_{u}}$, disturbances $w_{k}$ $\in$ $\mathcal{W}$ $\subseteq$ $\mathbb{R}^{n_{w}}$, such that
@@
We consider the problem of controlling this system between time step 0 and a known final time step NN $\in$ $\mathbb{N}$. We denote by $x\in\bar{\mathcal{X}}\coloneq\mathcal{X}^{N+1}$ the collection of states from $k=0$ to $k=N$, by $u\in\bar{\mathcal{U}}\coloneq\mathcal{U}^{N}$ the control inputs between 0 and N1N-1, and by $w\in\bar{\mathcal{W}}\coloneq\mathcal{W}^{N}$ the uncertainty sequence between 0 and N1N-1. The dynamics of (1) between 0 and $N$ can then be combined with an initial condition constraint on $x_{0}$ into a single function describing the evolution of the state,
@@
[…]
@@
In [16], we showed how the problem of (7) could be cast as a SIP and an open-loop solution could be obtained through the application of a local reduction method, as described in Section IV. Here, we examine the case where $u$ need not be computed in its entirety at time step 0, and show how updating $u$ within a decreasing horizon context can lead to performance improvements. Through the rest of this paper, we assume that the control inputs can be re-computed at every time step between 0 and $N$.
In order to describe our approach, we now introduce $x_{k}^{+}\in\mathcal{X}_{k}^{+}\coloneqq\mathcal{X}^{N-k}$ to refer to the components of $x$ corresponding to the time steps between $k+1$ and $N$, $u_{k}^{+}\in\mathcal{U}_{k}^{+}\coloneq\mathcal{U}^{N-k}$ to refer to the decisions of $u$ that are to be made between $k$ and N1N-1. Similarly, we define $w_{k}^{+}\in\mathcal{W}_{k}^{+}\coloneq\mathcal{W}^{N-k}$ to refer to the time-varying disturbances of $w$ that are realized between $k$ and $N$. We also use the notation $(\cdot)_{k}^{-}$ to refer to any quantity between 0 and k1k-1, with the exception of $x_{k}^{-}$, which includes $x_{0}$ through $x_{k}$.
@@
$J_{k}^{*}(z_{k},u_{k}^{+})\coloneqq\hskip-10.0pt\max_{\begin{subarray}{c}\vphantom{\sum}w_{k}^{+}\in\,\mathcal{W}_{k}^{+},\\[1.0pt] x_{k}^{+}\in\,\mathcal{X}_{k}^{+}(z_{k},u_{k}^{+},w_{k}^{+})\end{subarray}}\hskip-10.0ptJ_{k}(z_{k},x_{k}^{+},u_{k}^{+},w_{k}^{+}).$ (11)
@@
Meanwhile, if k<N1k<N-1 and we know that the controller will be updated in the future, we only need to compute the next segment in the controller trajectory $u^{*}_{k}(z_{k})$, the control input that, in the worst-case realization of $w_{k}$, leaves the system in the best possible position to minimize the total cost by the time the control input can next be updated at $k+1$. $u^{*}_{k}(z_{k})$ also needs to ensure that future constraints remain satisfiable for all possible uncertainties. The rest of the control trajectory, meanwhile, cannot be computed at time $k$ since it relies on information that will only be available in the future. In Section III, we show how to overcome this limitation to propose a novel update-aware robust optimal MPC algorithm that accounts for the acquisition of future information in its decision making process.
@@ III-A Algorithm Description
[…]
@@ III-B Theoretical Guarantees
Therefore, $\mathcal{U}_{k}^{*}(z_{k})$ and $\bar{\mathcal{U}}_{k}^{*}(z_{k})$ only differ by the ordering of the $\forall$ and $\exists$ operators. Then by universal instantiation, a member of $\bar{\mathcal{U}}_{k}^{*}(z_{k})$ must also be a member of ${\mathcal{U}}_{k}^{*}(z_{k})$, since $\exists A_{1}\forall A_{2}\implies\forall A_{2}\exists A_{1}$ given any $A_{1}$ and $A_{2}$. Members of ${\mathcal{U}}_{k}^{*}(z_{k})$, meanwhile, do not necessarily belong to $\bar{\mathcal{U}}_{k}^{*}(z_{k})$. ∎
@@
$\begin{split}&\bar{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})\coloneqq\left\{v_{k}\in\mathcal{U}\left|\,\exists v_{k+1}\in\mathcal{U}\left|\,\ldots\left|\,\exists v_{N}\in\mathcal{U}\left|\vphantom{\begin{array}[]{c}\\[1.0pt] \end{array}}\right.\right.\right.\right.\right.\\[1.0pt] &\left.\begin{array}[]{l}\forall{\omega}_{k}^{+}\in\mathcal{W}_{k}^{+},\\[1.0pt] \forall{s}_{k}^{+}\in X_{k}^{+}(z_{k},v_{k}^{+},\omega_{k}^{+})\end{array}\hskip-3.0pt:\hskip-3.0pt\right.\left.\>J_{k}(z_{k},s_{k}^{+},v_{k}^{+},\omega_{k}^{+})\leq\gamma_{k}\vphantom{\begin{array}[]{c}\\[1.0pt] \end{array}}\;\right\}.\end{split}$ (20)
Using the same logic as in Theorem 1, we know that $\bar{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})\subseteq{\mathcal{J}}^{*}_{k}(z_{k},\bar{\gamma_{k}})$ $\forall z_{k}\in\mathcal{Z}_{k},\gamma_{k}\in\mathbb{R}$. It then follows that $\left\{\bar{\mathcal{U}}_{k}^{*}(z_{k})\,\cap\,\bar{\mathcal{J}}^{*}(z_{k},\delta_{k})\right\}\subseteq\left\{{\mathcal{U}}_{k}^{*}(z_{k})\,\cap\,{\mathcal{J}}^{*}(z_{k},\delta_{k})\right\}$. Therefore, the feasible space for the problem of (19) is contained in that of (17b), and given that the two objectives are the same, we have that $\gamma_{k}(z_{1})\leq\bar{\gamma_{k}}(z_{k})$ $\forall z_{k}\in\mathcal{Z}_{k}$. ∎
@@
Proof:
Recursive feasibility follows trivially from (15) since $u^{*}_{k}(z_{k})\in\mathcal{U}_{k}^{*}(z_{k})\implies u_{k+1}^{*}(z_{k+1})\in\mathcal{U}_{k+1}^{*}(z_{k+1})$ $\forall z_{k}\in\mathcal{Z}_{k}$, $w_{k}\in\mathcal{W}$, where $u_{k}=u_{k}^{*}(z_{k})$, $x_{k}=X_{k}(z_{k},u_{k},w_{k})$, and $z_{k+1}=(z_{k},x_{k},u_{k},w_{k})$. ∎
@@
Proof:
This follows directly from the knowledge that $u_{k}\in\mathcal{U}^{*}(z_{k})\,\cap\,\mathcal{J}^{*}(z_{k},\gamma^{*}_{k})$ under (17a). After applying $u_{k}^{*}(z_{k})$, $\mathcal{J}_{k+1}^{*}(z_{k+1},\gamma_{k}^{*})$ must be non-empty from (16). Similarly, $\mathcal{U}^{*}_{k+1}(z_{k+1})$ is non-empty from Theorem 3. Therefore, $\gamma_{k}^{*}(z_{k})$ is guaranteed to be in the solution space of (17b) for $z_{k+1}$ and the optimal solution can only improve on the previous time-step so that $\gamma^{*}_{k+1}(z_{k+1})\leq\gamma^{*}_{k}(z_{k})$. ∎
@@ IV Solution by Local Reduction
While the problem of (17a) gives us the optimal continuation to our control input trajectory, it remains challenging to solve numerically. We address this by noticing that (17a) is a nested sequence of Nk+1N-k+1 existence-constrained robust optimal control problems, which, under Assumptions 14 can be rewritten as a nested sequence of existence-constrained SIPs using the approach of [16]. Then, from [16], (17) can be reformulated into the sequence of problems
@@
where $\tilde{P}_{k}(z_{k})\coloneqq P_{k}(z_{k})$ if $P_{k}(z_{k})$ is feasible and -$\infty$ if not, $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are members of a simplex $\Lambda$, and $\epsilon$ is the smallest positive constant. Substituting (27) for (24b) in (24) then finalizes the conversion into a regular SIP. Notice that solving the constraint generation problem for (27) requires solving $P_{k}^{k+1}(\cdot)$, preserving the recursive structure.
@@ V Results
In order to test the performance of our proposed update-aware robust optimal MPC algorithm, we compare it to a more standard robust optimal control scheme that applies the control of (12a) instead of (17a) on a simulation of a planar quadrotor. The model has continuous-time states $r,\dot{r},s,\dot{s},\psi,\dot{\psi}]$, where $r$ is the quadrotor’s horizontal position, $s$ is the quadrotor’s height, and $\psi$ is the quadrotor’s tilt angle, as illustrated in Figure 1.
@@
We then obtain the discrete-time dynamics at time step $k$ through the application of a midpoint rule with sampling time $T_{s}=0.5$, and where $x^{1}_{k}$ through $x^{6}_{k}$ are the discretized states at time $k$ corresponding to the continuous-time values $r,\dot{r},s,\dot{s},\psi,\dot{\psi}]$ and the discretization is such that $k=0$ maps to $t=0$. We examine the behavior of the simulation over 6 time steps, with $\mathcal{K}=\{0,1,2,3,4\}$.
The system is initialized from the origin with the state $x_{0}=[0,0,0,0,0,0]$. We also require that the system’s $x^{1}$ coordinate remain between c-c and $c$ at all times and that its $x^{3}$ coordinate remain positive by satisfying the inequalities
@@
[…]
@@
[…]